OpenStudy (anonymous):

A circular pond is modeled by the equipment x^2+y^2=225. A bridge over the pond is modeled by a segment of the equation x-7y=-75. What are the coordinates of the points where the bridge meets the edge of the pond?

OpenStudy (anonymous):

Solve simultaneously.

OpenStudy (anonymous):

I don't know what you mean

OpenStudy (anonymous):

For example \[x=7y-75\] then \[(7y-75)^2+y^2=225\] finally solve for the two values of y and substitute in the linear equation to get the corresponding x values

OpenStudy (anonymous):

the solutions should be (-12,9) and (9,12)

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

ok s I got till -50y^2-1050y+5400=0 But where do I go from there?

OpenStudy (anonymous):

Divide everything by -50

OpenStudy (anonymous):

Correction\[50*y^2-1050*y+5400\]

OpenStudy (anonymous):

\[y^2-21y+108=0\] implies\[(y-9)(y-12)=0\]

OpenStudy (anonymous):

How, did you factor or something?

OpenStudy (anonymous):

Yeah

OpenStudy (anonymous):

ok, also why did you divide by 50. Is it because it was in front of y^2

OpenStudy (anonymous):

We dont need the 50 since \[50(y^2-21y+108)=0\] iff \[y^2-21y+108=0\] by the zero product property

OpenStudy (anonymous):

oh ok, I get it. thanks